OpenStudy (christianm):

I don't know where to start on the problem 1a, from Unit 5, Integration Techniques. Could someone help me out?

OpenStudy (christianm):

My attempt is to draw a triangle, then evaluate \[\tan(\sqrt{3})\] then take the inverse, but this is not correct, since I would get \[\pi/6\] and an answer of \[6/\pi \] But the answer is \[\pi/3 \]